What is the principle of superposition of waves?
 

Grade:12

1 Answers

KALYAN
8 years ago
 Principle of superposition:
The principle of superposition of waves states that when two or more waves are simultaneously
interfere, the resultant displacement of any particle is equal to the algebraic sum of the
displacements of all the waves. If y1 = f1 ( x − vt ) ; y2 = f 2 ( x − vt ) ..... yn = f n ( x − vt ) Then
n
y = f1 ( x − vt ) + f 2 ( x − vt ) + ... + f n ( x − vt ) = ∑ f i ( x − vt )